### v1

@@ Abstract
dS/CFT, dS static holography, $T\bar{T}$ deformation.
@@ 1 Introduction
The holographic duality principle tHooft:1993dmi ; Susskind:1994vu offers a fundamental framework for studying quantum gravity. The duality between $(d+1)$-dimensional Anti-de Sitter (AdS) spacetime and $d$-dimensional conformal field theory (CFT) was well established Maldacena:1997re ; Gubser:1998bc ; Witten:1998qj . Recent developments have further extended this framework to include the $T\bar{T}$-deformed CFT Smirnov:2016lqw ; Cavaglia:2016oda , whose holographic dual was described by an AdS spacetime with a finite radial cutoff McGough:2016lol ; Kraus:2018xrn  [^3] Further interesting studies related to the $T\bar{T}$ deformation can be found in the reviews He:2025ppz ; Jiang:2019epa . . In contrast, the holographic description of de Sitter (dS) spacetime remains an open and intriguing area of research Witten:2001kn ; Chakraborty:2023yed ; Dey:2024zjx .
@@
In dS holography, two principal holographic frameworks have been proposed: the dS/CFT correspondence Strominger:2001pn ; Maldacena:2002vr ; Anninos:2011ui and the dS static patch holography Anninos:2011af ; Anninos:2017hhn ; Susskind:2021omt ; Susskind:2021esx ; Susskind:2021dfc ; Shaghoulian:2022fop ; Shaghoulian:2021cef ; Anninos:2024wpy ; Ruan:2025uhl . In the dS/CFT correspondence, the dual field theory is proposed to reside at the future spacelike asymptotic boundary, as inferred from the analysis of asymptotic symmetries Strominger:2001pn . In contrast, based on Bousso’s covariant entropy bound Bousso:1999xy , Refs. Susskind:2021omt ; Susskind:2021esx ; Susskind:2021dfc ; Shaghoulian:2022fop ; Shaghoulian:2021cef ; Anninos:2024wpy ; Ruan:2025uhl argued that the dual field theory should be located on the stretched horizon within the framework of dS static holography. Depending on the position of the holographic boundary, holographic models in dS spacetime can be broadly classified into two categories. The first class corresponds to scenarios where the dual field theory resides on a spacelike boundary, as in the dS/CFT correspondence and Cauchy slice holography Araujo-Regado:2022gvw ; Araujo-Regado:2022jpj , where the dual theory is nonunitary Hikida:2022ltr ; Araujo-Regado:2022gvw ; Doi:2022iyj ; Araujo-Regado:2022jpj . The second class introduces a timelike boundary in dS spacetime and places the dual field theory on this boundary, including worldline holography Anninos:2011af ; Anninos:2017hhn , dS static patch holography Susskind:2021omt , the dS/dS correspondence Alishahiha:2004md ; Alishahiha:2005dj ; Dong:2018cuv , and the half-dS holographic model Kawamoto:2023nki . In these cases, the dual field theory typically exhibits nonlocal features Miyaji:2015yva ; Geng:2019ruz ; Geng:2020kxh ; Kawamoto:2023nki ; Chang:2024voo . Moreover, a shift of the holographic boundary inward within the bulk induces different deformations on the field theory side. Specifically, the inward motion of a spacelike boundary corresponds to the $T\bar{T}$ deformation Chen:2023eic ; Chang:2024voo  [^4] In the $T\bar{T}$-deformed dS/CFT, both the central charge and the deformation parameter take complex values., whereas the inward motion of a timelike boundary corresponds to the $T\bar{T}+\Lambda_{2}$ deformation Gorbenko:2018oov ; Lewkowycz:2019xse ; Shyam:2021ciy ; Coleman:2021nor ; Torroba:2022jrk ; Silverstein:2022dfj ; Batra:2024kjl ; Aguilar-Gutierrez:2024nst .
@@
The origin of the $\Lambda_{2}$ term in the $T\bar{T}+\Lambda_{2}$ deformation has been attributed to a positive cosmological constant. In Ref. Gorbenko:2018oov , one starts with a $T\bar{T}$ flow; as the deformation parameter increases, the timelike dS2 boundary, on which the $T\bar{T}$-deformed CFT resides, moves from the AdS boundary toward the deep infrared region. When the deformation parameter approaches infinity, by exploiting the indistinguishability between AdS/dS and dS/dS geometries in the deep infrared region, the $\Lambda_{2}$ is turned on, and the $T\bar{T}+\Lambda_{2}$ flow emerges as the deformation parameter decreases. The resulting $T\bar{T}+\Lambda_{2}$ flow provides a mechanism to reconstruct the dS spacetime Gorbenko:2018oov . The $T\bar{T}+\Lambda_{2}$-deformed theory is conjectured to reside on a timelike dS2 boundary of dS3. In summary, the extended flow proposed above connects the two deformations listed in the second column of Table 1.
@@
In contrast to the construction joining AdS and dS throats discussed in Gorbenko:2018oov , the $T\bar{T}+\Lambda_{2}$ deformation can also arise in AdS spacetime with a spacelike boundary, as demonstrated by Ahmad, Almheiri, and Lin AliAhmad:2025kki . They employed a combined sequence of $T\bar{T}$ and $T\bar{T}+\Lambda_{2}$ deformations to probe the interior regions of a rotating BTZ black hole. In this setup, the $\Lambda_{2}$ term naturally emerges when the $T\bar{T}$ deformation pushes the spacetime boundary into the region between the outer and inner horizons, in which region, the boundary turns to spacelike while it is timelike outside the outer horizon. As the spacelike boundary is further pushed inward by the $T\bar{T}+\Lambda_{2}$ deformation, the $\Lambda_{2}$ term eventually vanishes once the boundary crosses the inner horizon and becomes timelike. Finally, the timelike boundary is then driven inward by the $T\bar{T}$ deformation all the way toward the black hole singularity. The two deformations that appear in the above analysis correspond to the two entries listed in the second row of Table 1.
@@
Motivated by Ref. AliAhmad:2025kki , we propose a unified framework for the two holographic models in dS holography by combining the $T\bar{T}$ and $T\bar{T}+\Lambda_{2}$ flows in dS spacetime, namely, relate two deformations in the third row of Table 1. Our construction is carried out in the static patch of dS spacetime and its extension. The static coordinate for dS spacetime is given by
where $r\in(0,\ell_{\rm dS})$ corresponds to the static region, whereas $r\in(\ell_{\rm dS},\infty)$ corresponds to its extended spacetime. The Penrose diagram of the dS static spacetime is shown in Fig. 1. In Fig. 1(a), $\mathcal{I}^{+}$ denotes the future infinity $r=\infty$, and $r=0$ represents the south pole, i.e., the position of a static observer. It is evident that outside the cosmological horizon ($r>\ell_{\rm dS}$), the coordinate $r$ behaves as timelike, while it remains spacelike inside the horizon. When the holographic principle is applied to the dS static spacetime, the holographic model outside the cosmological horizon corresponds to the first type of dS holographic model, whereas the model inside the horizon corresponds to the second type. In both cases, the dual field theory is defined on a codimension-one hypersurface at constant $r=r_{c}$.
@@
As illustrated in Fig. 1(b), when the boundary starts from spacelike infinity, it is driven inward by the $T\bar{T}$ deformation. Once the boundary crosses the cosmological horizon and becomes timelike, the deformation transitions to the $T\bar{T}+\Lambda_{2}$ one, which continues to push the boundary toward the static observer. This process offers a unified description of the two classes of holographic models in dS spacetime.
@@ 2 $T\bar{T}$ and $T\bar{T}+\Lambda_{2}$ deformations in dS spacetime
$H=\frac{16\pi G}{\sqrt{|h|}}\left[\Pi^{ab}\Pi_{ab}-({\Pi^{a}}_{a})^{2}\right]+n^{a}n_{a}\frac{\sqrt{|h|}}{16\pi G}\left(\tilde{\mathcal{R}}^{(2)}-2\Lambda\right)=0,$ (3)
where $\Lambda$ is the cosmological constant, given by $\Lambda=\frac{1}{\ell_{\rm dS}^{2}}$, and $n^{a}$ is the unit outward-pointing normal vector of the boundary at $r=r_{c}$, i.e., the direction towards increasing $r$. $\tilde{\mathcal{R}}^{(2)}$ denotes the intrinsic curvature of the boundary. And $\Pi^{ab}$ denotes the canonical momentum, which is defined as
@@
$n^{a}n_{a}\tilde{\mathcal{R}}^{(2)}+K^{2}-K_{ab}K^{ab}+2\Lambda n^{a}n_{a}=0.$ (5)
@@ 2.1 $T\bar{T}$ deformation with spacelike boundary
In the dS/CFT correspondence, since the CFT resides at future spacelike infinity with a Euclidean boundary metric, the boundary Brown–York tensor Brown:1992br ; Balasubramanian:1999re is defined as
$\tilde{T}_{ab}$ $=\frac{2}{\sqrt{h}}\frac{\delta}{\delta h^{ab}}(-\ln Z_{\text{dS}})$ (9)
@@
$=\frac{i}{8\pi G}\left(K_{ab}-Kh_{ab}+\frac{h_{ab}}{\ell_{\rm dS}}\right).$
For a spacelike boundary with $n^{a}n_{a}=-1$, the Hamiltonian constraint (5) reduces to the standard $2+1$ decomposition form,
$\tilde{\mathcal{R}}^{(2)}+K^{2}-K_{ab}K^{ab}-2\Lambda=0.$ (10)
Rewriting the extrinsic curvature in terms of the boundary Brown–York tensor yields the bulk trace flow equation,
$\tilde{T}^{a}_{a}=\frac{i\ell_{\rm dS}}{16\pi G}\tilde{\mathcal{R}}^{(2)}+i4\pi G\ell_{\rm dS}\left(\tilde{T}^{ab}\tilde{T}_{ab}-(\tilde{T}^{a}_{a})^{2}\right).$ (11)
Comparing this with the trace flow equation in the dual field theory,
$T^{a}_{a}=-\frac{c}{24\pi}\mathcal{R}^{(2)}-4\pi\lambda\,\langle T\bar{T}\rangle,\quad T\bar{T}\equiv\frac{1}{8}\left(T^{ab}T_{ab}-(T^{a}_{a})^{2}\right),$ (12)
@@
It should be noted that the boundary Brown–York tensor and the stress tensor of the dual CFT differ by a conformal factor depending on the radial coordinate $r$. Near spacelike infinity, the static coordinates of dS spacetime can be written in the Fefferman–Graham form,
@@
$\tilde{\mathcal{R}}^{(2)}=\frac{1}{r^{2}}\mathcal{R}^{(2)},\qquad\tilde{T}_{ab}=T_{ab},\qquad\tilde{T}^{ab}=\frac{1}{r^{4}}T^{ab},\qquad\tilde{T}^{a}_{a}=\frac{1}{r^{2}}T^{a}_{a}.$ (15)
@@
$T=-\frac{\pi\lambda}{2}\left(T^{ab}T_{ab}-(T^{a}_{a})^{2}\right).$ (16)
@@
$\delta(-\ln Z_{\text{QFT}})=\frac{1}{2}\int d^{2}x\sqrt{\gamma}\,T_{ab}\delta\gamma^{ab}=\frac{1}{2\lambda}\int d^{2}x\sqrt{\gamma}\,T\,\delta\lambda.$ (19)
@@
$\partial_{\lambda}\ln Z_{\text{QFT}}=-\frac{1}{2\lambda}\int d^{2}x\sqrt{\gamma}\,T=\frac{\pi}{4}\int d^{2}x\sqrt{\gamma}\left(T^{ab}T_{ab}-(T^{a}_{a})^{2}\right).$ (20)
@@
$\partial_{\lambda_{\rm dS}}S_{\text{QFT}}=\frac{\pi}{4}\int d^{2}x\,\sqrt{\gamma}\,\left(T^{ab}T_{ab}-(T^{a}_{a})^{2}\right).$ (21)

### v2

@@ Abstract
Holographic duality, dS holography, $T\bar{T}$ deformation.
@@ 1 Introduction
The holographic duality principle tHooft:1993dmi ; Susskind:1994vu offers a fundamental framework for studying quantum gravity. The duality between $(d+1)$-dimensional Anti-de Sitter (AdS) spacetime and $d$-dimensional conformal field theory (CFT) was well established Maldacena:1997re ; Gubser:1998bc ; Witten:1998qj . Recent developments have further extended this framework to include the $T\bar{T}$-deformed CFT Smirnov:2016lqw ; Cavaglia:2016oda , whose holographic dual was described by an AdS spacetime with a finite radial cutoff McGough:2016lol ; Kraus:2018xrn  [^3] Further interesting studies related to the $T\bar{T}$ deformation can be found in the reviews He:2025ppz ; Jiang:2019epa . . In contrast, the holographic description of de Sitter (dS) spacetime remains an open and intriguing area of research Witten:2001kn ; Chakraborty:2023yed ; Dey:2024zjx ; Yadav:2024ray ; Goodhew:2024eup ; Thavanesan:2025kyc ; Chakravarty:2025sbg .
@@
In dS holography, two principal holographic frameworks have been proposed: the dS/CFT correspondence Strominger:2001pn ; Maldacena:2002vr ; Anninos:2011ui and the dS static patch holography Anninos:2011af ; Anninos:2017hhn ; Susskind:2021omt ; Susskind:2021esx ; Susskind:2021dfc ; Shaghoulian:2022fop ; Shaghoulian:2021cef ; Franken:2023pni ; Franken:2023jas ; Anninos:2024wpy ; Ruan:2025uhl . In the dS/CFT correspondence, the dual field theory is proposed to reside at the future spacelike asymptotic boundary, as inferred from the analysis of asymptotic symmetries Strominger:2001pn . In contrast, based on Bousso’s covariant entropy bound Bousso:1999xy , Refs. Susskind:2021omt ; Susskind:2021esx ; Susskind:2021dfc ; Shaghoulian:2022fop ; Shaghoulian:2021cef ; Anninos:2024wpy ; Ruan:2025uhl argued that the dual field theory should be located on the stretched horizon within the framework of dS static holography. Depending on the position of the holographic boundary, holographic models in dS spacetime can be broadly classified into two categories. The first class corresponds to scenarios where the dual field theory resides on a spacelike boundary, as in the dS/CFT correspondence and Cauchy slice holography Araujo-Regado:2022gvw ; Araujo-Regado:2022jpj ; Araujo-Regado:2025elv , where the dual theory is nonunitary Hikida:2022ltr ; Araujo-Regado:2022gvw ; Doi:2022iyj ; Araujo-Regado:2022jpj . The second class introduces a timelike boundary in dS spacetime and places the dual field theory on this boundary, including worldline holography Anninos:2011af ; Anninos:2017hhn , dS static patch holography Susskind:2021omt , the dS/dS correspondence Alishahiha:2004md ; Alishahiha:2005dj ; Dong:2018cuv ; Geng:2019bnn ; Grieninger:2019zts , and the half-dS holographic model Kawamoto:2023nki . In these cases, the dual field theory typically exhibits nonlocal features Miyaji:2015yva ; Geng:2019ruz ; Geng:2020kxh ; Kawamoto:2023nki ; Chang:2024voo . Moreover, a shift of the holographic boundary inward within the bulk induces different deformations on the field theory side. Specifically, the inward motion of a spacelike boundary corresponds to the $T\bar{T}$ deformation Chen:2023eic ; Chang:2024voo , whereas the inward motion of a timelike boundary corresponds to the $T\bar{T}+\Lambda_{2}$ deformation Gorbenko:2018oov ; Lewkowycz:2019xse ; Shyam:2021ciy ; Coleman:2021nor ; Torroba:2022jrk ; Silverstein:2022dfj ; Batra:2024kjl ; Aguilar-Gutierrez:2024nst .
Following the proposal of the $T\bar{T}$ deformation in the context of the AdS/CFT correspondence, deformations of dS holographic models within the dS/dS correspondence have also been investigated. It was proposed that, in dS holography, shifting the timelike boundary inward into the bulk is dual to the $T\bar{T}+\Lambda_{2}$ deformation on the field theory side, with the $\Lambda_{2}$ term originating from the positive cosmological constant of dS spacetime Gorbenko:2018oov . This proposal is supported, in particular, by the trace flow equation
$\tilde{T}^{i}_{i}=-\frac{\ell_{\text{(A)dS}}}{16\pi G}\tilde{\mathcal{R}}^{(2)}-4\pi G\ell_{\text{(A)dS}}\left(\tilde{T}^{ij}\tilde{T}_{ij}-(\tilde{T}^{i}_{i})^{2}\right)-\frac{\eta-1}{8\pi G\ell_{\text{(A)dS}}}\,$, (1)
where $\tilde{T}_{ij}$ and $\tilde{\mathcal{R}}^{(2)}$ denote the boundary Brown–York tensor and the intrinsic curvature of the timelike boundary, respectively, while $\ell_{\text{(A)dS}}$ represents the curvature radius of (A)dS spacetime. The parameter $\eta$ is related to the cosmological constant through $\Lambda_{2}=-\eta/\ell_{\text{(A)dS}}^{2}$. For AdS spacetime one has $\eta=1$ and $\Lambda_{2}=-1/(\ell_{\text{AdS}})^{2}$, whereas for dS spacetime $\eta=-1$ and $\Lambda_{2}=1/\ell_{\text{dS}}^{2}$. This implies that, upon identifying the corresponding bulk and boundary quantities, the trace flow equation on the dual field theory side of dS spacetime takes the form
$T^{i}_{i}=-\frac{c}{24\pi}\mathcal{R}^{(2)}-4\pi\lambda\,\langle T\bar{T}\rangle+\frac{2}{\pi\lambda}\,.$ (2)
Here $T_{ij}$ and $\mathcal{R}^{(2)}$ denote the stress tensor of the boundary field theory and the intrinsic curvature of the manifold on which the field theory is defined, respectively, and they differ from the corresponding bulk quantities $\tilde{T}_{ij}$ and $\tilde{\mathcal{R}}^{(2)}$ by Weyl factors, since they do not depend on the radial coordinate. Compared to the standard $T\bar{T}$ deformation, this expression contains an additional term associated with the positive cosmological constant. As a direct consequence of the trace flow equation, the partition function of the dual field theory satisfies the flow equation
$\frac{\partial}{\partial\lambda}\log Z=-2\pi\int d^{2}x\sqrt{g}\langle T\bar{T}\rangle+\frac{1}{\pi\lambda^{2}}\int d^{2}x\sqrt{g}\,.$ (3)
This proposal was further supported by consistency checks based on holographic entanglement entropy Lewkowycz:2019xse . Moreover, within the framework of dS static patch holography, the entropy of the stretched horizon and its logarithmic corrections were computed by employing the $T\bar{T}+\Lambda_{2}$ deformation Shyam:2021ciy .
Subsequently, deformations of the dS/CFT correspondence were explored in Ref. Chen:2023eic . It was proposed that the $T\bar{T}$ deformation of the dual field theory is holographically dual to moving the spacelike boundary at future infinity inward into the bulk. In contrast to the AdS/CFT case, both the deformation parameter and the central charge in the dS/CFT setup are purely imaginary. It was further shown in Ref. Chang:2024voo that the trace flow equation in dS/CFT can be obtained via analytic continuation of the corresponding trace flow equation in AdS/CFT through a double Wick rotation, which naturally renders both quantities imaginary. To test the proposed holographic duality of the deformed field theory, Ref. Chen:2023eic employed the sphere partition function to compute exactly the entanglement entropy of a meridian on the boundary sphere. These results were subsequently generalized in Ref. Chang:2024voo to a variety of dS backgrounds, including dS global spacetime and the dS Poincaré patch, where consistency between holographic entanglement entropy and geodesic lengths was examined.
@@
Composite flows formed by combining the $T\bar{T}$ flow with the $T\bar{T}+\Lambda_{2}$ flow have found wide-ranging applications. In Ref. Gorbenko:2018oov , the construction begins with a $T\bar{T}$ flow; as the deformation parameter increases, the timelike dS2 boundary, on which the $T\bar{T}$-deformed CFT resides, moves from the AdS boundary toward the deep infrared. In this limit, by exploiting the indistinguishability of the AdS/dS and dS/dS geometries in the deep infrared regime, the $\Lambda_{2}$ term is introduced, and the $T\bar{T}+\Lambda_{2}$ flow emerges as the deformation parameter decreases. In essence, such a composite flow provides a mechanism for reconstructing dS spacetime. Moreover, Ref. Coleman:2021nor similarly exploited the near-horizon indistinguishability between the black hole horizon and the cosmological horizon to perform a microscopic counting of states in dS spacetime, thereby providing a microscopic description of the cosmological horizon entropy. In summary, the extended flow proposed above connects the two deformations listed in the second column of Table 1.
@@
In contrast to the construction joining AdS and dS throats discussed in Gorbenko:2018oov , the $T\bar{T}+\Lambda_{2}$ deformation can also arise in AdS spacetime with a spacelike boundary, as demonstrated by Ahmad, Almheiri, and Lin AliAhmad:2025kki . They employed a combined sequence of $T\bar{T}$ and $T\bar{T}+\Lambda_{2}$ deformations to probe the interior regions of a rotating BTZ black hole. In this setup, the $\Lambda_{2}$ term naturally emerges when the $T\bar{T}$ deformation pushes the spacetime boundary into the region between the outer and inner horizons, in which region, the boundary turns to spacelike while it is timelike outside the outer horizon. As the spacelike boundary is further pushed inward by the $T\bar{T}+\Lambda_{2}$ deformation, the $\Lambda_{2}$ term eventually vanishes once the boundary crosses the inner horizon and becomes timelike. Finally, the timelike boundary is then driven inward by the $T\bar{T}$ deformation all the way toward the black hole singularity. The two deformations that appear in the above analysis correspond to the two entries listed in the second row of Table 1444The $T\bar{T}+\Lambda_{2}$ deformation denotes a class of standard $T\bar{T}$ deformation plus a constant term, and is not restricted to dS spacetime..
@@
As discussed above, the two classes of models in dS holography exhibit pronounced differences in the fundamental properties of their dual field theories. Holography with a spacelike boundary is typically associated with a non-unitary dual field theory, whereas holography with a timelike boundary often gives rise to a dual theory that is intrinsically nonlocal. This inherent tension between unitarity and locality constitutes one of the central features that distinguish dS holography from the AdS/CFT correspondence, and it highlights the fundamental challenges in constructing a consistent holographic dual description of dS spacetime. Motivated by Ref. AliAhmad:2025kki , we propose a consistent dS holographic model by combining the $T\bar{T}$ and $T\bar{T}+\Lambda_{2}$ flows in dS spacetime, as summarized in the third row of Table 1. This construction provides a unified framework that simultaneously incorporates both classes of dS holographic models.
@@
Our construction is carried out in the static patch of dS spacetime and its extension. The static coordinate for dS spacetime is given by
where $r\in(0,\ell_{\rm dS})$ corresponds to the static region, whereas $r\in(\ell_{\rm dS},\infty)$ corresponds to its extended spacetime. The Penrose diagram of the dS static spacetime is shown in Fig. 1. In Fig. 1(a), $\mathcal{I}^{+}$ denotes the future infinity $r=\infty$, and $r=0$ represents the south pole, i.e., the position of a static observer. It is evident that outside the cosmological horizon ($r>\ell_{\rm dS}$), the coordinate $r$ behaves as timelike, while it remains spacelike inside the horizon. When the holographic principle is applied to the dS static spacetime, the holographic model outside the cosmological horizon corresponds to the first type of dS holographic model, whereas the model inside the horizon corresponds to the second type. In both cases, the dual field theory is defined on a codimension-one hypersurface at constant $r=r_{c}$. As illustrated in Fig. 1(b), when the boundary starts from spacelike infinity, it is driven inward by the $T\bar{T}$ deformation. Once the boundary crosses the cosmological horizon and becomes timelike, the deformation transitions to the $T\bar{T}+\Lambda_{2}$ one, which continues to push the boundary toward the static observer. This process offers a unified description of the two classes of holographic models in dS spacetime.
@@ 2 $T\bar{T}$ and $T\bar{T}+\Lambda_{2}$ deformations in dS spacetime
$H=\frac{16\pi G}{\sqrt{|h|}}\left[\Pi^{ab}\Pi_{ab}-({\Pi^{a}}_{a})^{2}\right]+n^{a}n_{a}\frac{\sqrt{|h|}}{16\pi G}\left(\tilde{\mathcal{R}}^{(2)}-2\Lambda_{2}\right)=0,$ (6)
where $\Lambda_{2}$ is the cosmological constant, given by $\Lambda_{2}=\frac{1}{\ell_{\rm dS}^{2}}$, and $n^{a}$ is the unit outward-pointing normal vector of the boundary at $r=r_{c}$, i.e., the direction towards increasing $r$. $\tilde{\mathcal{R}}^{(2)}$ denotes the intrinsic curvature of the boundary. And $\Pi^{ab}$ denotes the canonical momentum, which is defined as
@@
$n^{a}n_{a}\tilde{\mathcal{R}}^{(2)}+K^{2}-K_{ab}K^{ab}+2\Lambda_{2}n^{a}n_{a}=0.$ (8)
@@ 2.1 $T\bar{T}$ deformation with spacelike boundary
In the dS/CFT correspondence, since the CFT resides at future spacelike infinity with a Euclidean boundary metric, the Brown–York tensor of the Euclidean boundary Brown:1992br ; Balasubramanian:1999re is defined as
$\widetilde{T}_{(E)\,ab}$ $=\frac{2}{\sqrt{h}}\frac{\delta}{\delta h^{ab}}(-\ln Z_{\text{dS}})$ (12)
@@
$=\frac{i}{8\pi G}\left(K_{ab}-Kh_{ab}+\frac{h_{ab}}{\ell_{\rm dS}}\right),$
where the subscript $(E)$ denotes that the stress tensor is defined in Euclidean signature, while $(L)$ for Lorentzian signature in the next subsection. For a spacelike boundary with $n^{a}n_{a}=-1$, the Hamiltonian constraint (8) reduces to the standard $2+1$ decomposition form,
$\widetilde{\mathcal{R}}^{(2)}+K^{2}-K_{ab}K^{ab}-2\Lambda_{2}=0.$ (13)
Rewriting the extrinsic curvature in terms of the Brown–York tensor yields the bulk trace flow equation,
$\widetilde{T}_{(E)}=\frac{i\ell_{\rm dS}}{16\pi G}\widetilde{\mathcal{R}}^{(2)}+i4\pi G\ell_{\rm dS}\left(\widetilde{T}_{(E)}^{ab}\widetilde{T}_{(E)\,ab}-(\widetilde{T}_{(E)})^{2}\right).$ (14)
Comparing this with the trace flow equation in the dual field theory  [^6] Through out this work, the tilde quantity $\widetilde{T}_{ab}$ and untilde one $T_{ab}$ denote the stress tensor computed from bulk and boundary field theory, respectively.,
$T_{(E)}=-\frac{c}{24\pi}\mathcal{R}^{(2)}-4\pi\lambda\,\langle T\bar{T}\rangle,\quad T\bar{T}\equiv\frac{1}{8}\left(T_{(E)}^{ab}T_{(E)\,ab}-T_{(E)}^{2}\right),$ (15)
@@
It should be noted that the Brown–York tensor and the stress tensor of the dual CFT differ by a conformal factor depending on the radial coordinate $r$. Near spacelike infinity, the static coordinates of dS spacetime can be written in the Fefferman–Graham form,
@@
$\widetilde{\mathcal{R}}^{(2)}=\frac{1}{r^{2}}\mathcal{R}^{(2)},\;\widetilde{T}_{(E)\,ab}=T_{(E)\,ab},\;\widetilde{T}_{(E)}^{ab}=\frac{1}{r^{4}}T_{(E)}^{ab},\;\widetilde{T}_{(E)}=\frac{1}{r^{2}}T_{(E)}.$ (18)
@@
$T_{(E)}=-\frac{\pi\lambda}{2}\left(T_{(E)}^{ab}T_{(E)\,ab}-T_{(E)}^{2}\right).$ (19)
@@
$\delta(-\ln Z_{\text{QFT}})=\frac{1}{2}\int d^{2}x\sqrt{\gamma}\,T_{(E)\,ab}\delta\gamma^{ab}=\frac{1}{2\lambda}\int d^{2}x\sqrt{\gamma}\,T_{(E)}\,\delta\lambda.$ (22)
@@
$\partial_{\lambda}\ln Z_{\text{QFT}}=-\frac{1}{2\lambda}\int d^{2}x\sqrt{\gamma}\,T_{(E)}=\frac{\pi}{4}\int d^{2}x\sqrt{\gamma}\left(T_{(E)}^{ab}T_{(E)\,ab}-T_{(E)}^{2}\right).$ (23)
@@
$\partial_{\lambda_{\rm dS}}S_{\text{QFT}}=\frac{\pi}{4}\int d^{2}x\,\sqrt{\gamma}\,\left(T_{(E)}^{ab}T_{(E)\,ab}-T_{(E)}^{2}\right).$ (24)